# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ I.3 Over-specified Setting
In this appendix, we employ results for M-estimators in [89] to establish the density estimation rate under the top-K sparse Laplace gating Gaussian mixture of experts (MoE).
[…]
Lemma K.1.
If $\Omega$ is a bounded set, then the following inequalities hold for any $20<\eta<1/$ 2:
  • (i)
    $\log N(\eta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta)$;
  • (ii)
    $H_{B}(\eta,\mathcal{P}_{k_{*}}(\Omega),h)\lesssim\log(1/\eta)$.
Proof of Lemma K.1 is in Appendix K.1.2. Subsequently, we denote
$\widetilde{\mathcal{P}}_{k_{*}}(\Omega)$ :=$:=\{p_{(G+G_{*})/2}(Y|X):G\in\mathcal{G}_{k_{*}}(\Omega)\};$
$\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega)$ :=$:=\{p^{1/2}_{(G+G_{*})/2}(Y|X):G\in\mathcal{G}_{k_{*}}(\Omega)\}.$
In addition, for each $\delta>0$, we define a Hellinger ball centered around the conditional density function $p_{G_{*}}(Y|X)$ and intersected with the set $\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega)$ as
$\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,\delta):=\{p^{1/2}\in\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega):h(p,p_{G_{*}})\leq\delta\}.$
To capture the size of the above Hellinger ball, [89] suggest using the following quantity:
$\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,\delta)):=\int_{\delta^{2}/2^{13}}^{\delta}H^{1/2}_{B}(t,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,t),\|\cdot\|)\mathrm{d}t\vee\delta,$ (11)
where $t\vee\delta:=\max\{t,\delta\}$. Given those notations, let us recall a standard result for density estimation in [89].
Lemma K.2 (Theorem 7.4, [89]).
Take $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,\delta))$ such that $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Then, for some sequence $(\delta_{n})$ and universal constant $c$ which satisfy $\sqrt{n}\delta^{2}_{n}\geq c\Psi(\delta)$, we obtain that
for any $\delta\geq\delta_{n}$
Proof of Lemma K.2 can be found in [89]. Now, we are ready to provide the proof for convergence rate of density estimation in Theorem J.1 in Appendix K.1.1.
K.1.1 Main Proof
It is worth noting that for any $t>0$, we have
$H_{B}(t,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,t),\|\cdot\|)\leq H_{B}(t,\mathcal{P}_{k_{*}}(\Omega,t),h).$
Then, the integral in equation 11 is upper bounded as follows:
$\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,\delta))\leq\int_{\delta^{2}/2^{13}}^{\delta}H^{1/2}_{B}(t,\mathcal{P}_{k_{*}}(\Omega,t),h)\mathrm{d}t\vee\delta\lesssim\int_{\delta^{2}/2^{13}}^{\delta}\log(1/t)\mathrm{d}t\vee\delta,$ (12)
where the second inequality follows from part (ii) of Lemma K.1.
As a result, by choosing $\Psi(\delta)=\delta\cdot\sqrt{\log(1/\delta)}$, we can verify that $\Psi(\delta)/\delta^{2}$ is a non-increasing function of $\delta$. Furthermore, the inequality in equation 12 indicates that $\Psi(\delta)\geq\mathcal{J}_{B}(\delta,\widetilde{\mathcal{P}}^{1/2}_{k_{*}}(\Omega,\delta))$. Next, let us consider a sequence $(\delta_{n})$ defined as $\delta_{n}:=\sqrt{\log(n)/n}$. This sequence can be validated to satisfy the condition $\sqrt{n}\delta_{n}^{2}\geq c\Psi(\delta)$ for some universal constant $c$. Therefore, by Lemma K.2, we reach the conclusion of Theorem J.1:
for some universal constant $C$ depending only on $\Omega$.
K.1.2 Proof of Lemma K.1
Part (i). In this part, we will derive the following upper bound for the covering number of metric space $(\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})$ for any $20<\eta<1/$ 2 given the bounded set $\Omega$:
$\log N(\eta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta).$
To start with, we denote $\Theta:=\{(a,b,\nu)\in\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}_{+}:(\beta,W,a,b,\nu)\in\Omega\}$. As $\Omega$ is a bounded set, the set $\Theta$ is also bounded. Therefore, we can find an $\eta$-cover of $\Theta$, denoted by $\overline{\Theta}_{\eta}$. Additionally, we also define $\Delta:=\{(\beta,W)\in\mathbb{R}\times\mathbb{R}^{d}:(\beta,W,a,b,\nu)\in\Omega\}$, and $\overline{\Delta}_{\eta}$ be an $\eta$-cover of $\Delta$. Then, it can be validated that
$|\overline{\Theta}_{\eta}|\leq\mathcal{O}(\eta^{-(d+2)k_{*}}),\quad|\overline{\Delta}_{\eta}|\leq\mathcal{O}(\eta^{-(d+1)k_{*}}).$
Next, for each mixing measure $G=\sum_{i=1}^{k_{*}}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}\in\mathcal{G}_{k_{*}}(\Omega)$, we take into account two other mixing measures. The first measure is $G^{\prime}=\sum_{i=1}^{k_{*}}\exp(\beta_{i})\delta_{(W_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}$, where $(\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Theta}_{\eta}$ is the closest points to $(a_{i},b_{i},\nu_{i})$ in this set for all $i\in[k_{*}]$. The second one is $\overline{G}:=\sum_{i=1}^{k_{*}}\exp(\overline{\beta}_{i})\delta_{(\overline{W}_{i},\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})}$ in which $(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta}$ for any $i\in[k_{*}]$. Next, let us define
$\mathcal{T}:=\{p_{\overline{G}}\in\mathcal{P}_{k_{*}}(\Omega):(\overline{\beta}_{i},\overline{W}_{i})\in\overline{\Delta}_{\eta},\ (\overline{a}_{i},\overline{b}_{i},\overline{\nu}_{i})\in\overline{\Theta}_{\eta},\forall i\in[k_{*}]\},$
then it is obvious that $p_{\overline{G}}\in\mathcal{T}$. Now, we will show that $\mathcal{T}$ is an $\eta$-cover of metric space $(\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})$ with a note that it is not necessarily the smallest cover. Indeed, according to the triangle inequality,
$\|p_{G}-p_{\overline{G}}\|_{1}\leq\|p_{G}-p_{G^{\prime}}\|_{1}+\|p_{G^{\prime}}-p_{\overline{G}}\|_{1}.$ (13)
$\|p_{G}-p_{G^{\prime}}\|_{1}$ $\leq\sum_{i=1}^{k_{*}}\int\Big{|}f(Y|a_{i}^{\top}X+b_{i},\nu_{i})-f(Y|\overline{a}_{i}^{\top}X+\overline{b}_{i},\overline{\nu}_{i})\Big{|}\mathrm{d}(X,Y)$
$\lesssim\sum_{i=1}^{k_{*}}\int\Big{(}\|a_{i}-\overline{a}_{i}\|+\|b_{i}-\overline{b}_{i}\|+\|\nu_{i}-\overline{\nu}_{i}\|\Big{)}\mathrm{d}(X,Y)$
$=\sum_{i=1}^{k_{*}}\Big{(}\|a_{i}-\overline{a}_{i}\|+\|b_{i}-\overline{b}_{i}\|+\|\nu_{i}-\overline{\nu}_{i}\|\Big{)}$
$\lesssim\eta.$ (14)
Subsequently, we show that $\|p_{G^{\prime}}-p_{\overline{G}}\|_{1}\lesssim\eta$. For that purpose, we consider $q:=\binom{k_{*}}{K}$ $K$-element subsets of $\{1,\ldots,k_{*}\}$, which are assumed to take the form $\{\tau_{1},\tau_{2},\ldots,\tau_{K}\}$ for any $\tau\in[q]$. Additionally, we also denote $\{\tau_{K+1},\ldots,\tau_{k_{*}}\}:=\{1,\ldots,k_{*}\}\setminus\{\tau_{1},\ldots,\tau_{K}\}$ for any $\tau\in[q]$. Then, we define
$\mathcal{X}_{\tau}$ :=$:=\{x\in\mathcal{X}:-\|W_{i}-x\|\geq-\|W_{i^{\prime}}-x\|:i\in\{\tau_{1},\ldots,\tau_{K}\},i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\},$
$\widetilde{\mathcal{X}}_{\tau}$ :=$:=\{x\in\mathcal{X}:-\|\overline{W}_{i}-x\|\geq-\|\overline{W}_{i^{\prime}}-x\|:i\in\{\tau_{1},\ldots,\tau_{K}\},i^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\}.$
By using the same arguments as in the proof of Lemma L.1 in Appendix L, we achieve that either $\mathcal{X}_{\tau}=\widetilde{\mathcal{X}}_{\tau}$ or $\mathcal{X}_{\tau}$ has measure zero for any $\tau\in[q]$. As the $\mathrm{softmax}$ function is differentiable, it is a Lipschitz function with some Lipschitz constant $L\geq 0$. Next, we denote
$\pi_{{\tau}}(X):=\Big{(}-\|W_{\tau_{i}}-X\|+\beta_{{\tau}_{i}}\Big{)}_{i=1}^{{K}};\qquad\overline{\pi}_{{\tau}}(X):=\Big{(}-\|\overline{W}_{{\tau}_{i}}-X\|+\overline{\beta}_{{\tau}_{i}}\Big{)}_{i=1}^{{K}},$
for any ${K}$-element subset $\{{\tau}_{1},\ldots{\tau}_{K}\}$ of $\{1,\ldots,k_{*}\}$. Then, we get
$\|\mathrm{softmax}(\pi_{{\tau}}(X))-\mathrm{softmax}(\overline{\pi}_{{\tau}}(X))\|$ $\leq L\cdot\|\pi_{{\tau}}(X)-\overline{\pi}_{{\tau}}(X)\|$
$\leq L\cdot\sum_{i=1}^{{K}}\Big{(}\Big{|}\|W_{{\tau}_{i}}-X\|-\|\overline{W}_{{\tau}_{i}}-X\|\Big{|}+|\beta_{{\tau}_{i}}-\overline{\beta}_{{\tau}_{i}}|\Big{)}$
$\leq L\cdot\sum_{i=1}^{{K}}\Big{(}\|W_{{\tau}_{i}}-\overline{W}_{{\tau}_{i}}\|+|\beta_{{\tau}_{i}}-\overline{\beta}_{{\tau}_{i}}|\Big{)}$
$\leq L\cdot\sum_{i=1}^{{K}}\Big{(}\eta+\eta\Big{)}$
$\lesssim\eta.$
Back to the proof for $\|p_{G^{\prime}}-p_{\overline{G}}\|_{1}\lesssim\eta$, it follows from the above results that
$\|p_{G^{\prime}}-p_{\overline{G}}\|_{1}$ $=\int|p_{G^{\prime}}(Y|X)-p_{\overline{G}}(Y|X)|~{}\mathrm{d}(X,Y)$
$\leq\sum_{\tau=1}^{q}\int_{\mathcal{X}_{\tau}\times\mathcal{Y}}|p_{G^{\prime}}(Y|X)-p_{\overline{G}}(Y|X)|~{}\mathrm{d}(X,Y)$
$\leq\sum_{\tau=1}^{q}\int_{\mathcal{X}_{\tau}\times\mathcal{Y}}\sum_{i=1}^{K}\Big{|}\mathrm{softmax}(\pi_{\tau}(X)_{i})-\mathrm{softmax}(\overline{\pi}_{\tau}(X)_{i})\Big{|}\cdot\Big{|}f(Y|\overline{a}_{\tau_{i}}^{\top}X+\overline{b}_{\tau_{i}},\overline{\nu}_{\tau_{i}})\Big{|}~{}\mathrm{d}(X,Y)$
$\lesssim\eta,$ (15)
It follows from the results in equation 13, equation 14 and equation 15 that $\|p_{G}-p_{\overline{G}}\|_{1}\lesssim\eta$. This result indicates that $\mathcal{T}$ is an $\eta$-cover of the metric space $(\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})$. As a consequence, we obtain that
$N(\eta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\lesssim|\overline{\Delta}_{\eta}|\times|\overline{\Theta}_{\eta}|\leq\mathcal{O}(1/\eta^{(2d+3)k}),$
which leads to the conclusion of this part: $\log N(\eta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta)$.
Part (ii). In this part, we provide an upper bound for the bracketing entropy of $\mathcal{P}_{k_{*}}(\Omega)$ under the Hellinger distance $h$:
$H_{B}(\eta,\mathcal{P}_{k_{*}}(\Omega),h)\lesssim\log(1/\eta).$
Since $\Omega$ and $\mathcal{X}$ are bounded sets, there exist positive constants $\gamma,\ell,u$ such that $-\gamma\leq a^{\top}X+b\leq\gamma$ and $\ell\leq\nu\leq u$. Let us define
Next, let $\zeta\leq\eta$ which will be chosen later and $\{p_{1},\ldots,p_{N}\}$ be an $\zeta$-cover of metric space $(\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})$ with the covering number $N:=N(\tau,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})$. Additionally, we also consider brackets of the form $[\Psi_{i}^{L}(Y|X),\Psi_{i}^{U}(Y|X)]$ where
$\Psi_{i}^{L}(Y|X)$ $:=\max\{p_{i}(Y|X)-\tau,0\}$
$\Psi_{i}^{U}(Y|X)$ $:=\max\{p_{i}(Y|X)+\tau,B(Y|X)\}.$
Then, we can check that $\mathcal{P}_{k_{*}}(\Omega)\subseteq\bigcup_{i=1}^{N}[\Psi_{i}^{L}(Y|X),\Psi_{i}^{U}(Y|X)]$ and $\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)\leq\min\{2\eta,B(Y|X)\}$.
$\|\Psi_{i}^{U}-\Psi_{i}^{L}\|_{1}$ $=\int_{|Y|<2\gamma}[\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)]~{}\mathrm{d}X\mathrm{d}Y+\int_{|Y|\geq 2\gamma}[\Psi_{i}^{U}(Y|X)-\Psi_{i}^{L}(Y|X)]~{}\mathrm{d}X\mathrm{d}Y$
$\leq S\zeta+\exp\Big{(}-\frac{S^{2}}{2u}\Big{)}\leq S^{\prime}\zeta,$
$H_{B}(S^{\prime}\zeta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\leq\log N(\zeta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\leq\log(1/\tau).$
By setting $\zeta=\eta/S^{\prime}$, we obtain that $H_{B}(\eta,\mathcal{P}_{k_{*}}(\Omega),\|\cdot\|_{1})\lesssim\log(1/\eta)$. Finally, since the norm $\|\cdot\|_{1}$ is upper bounded by the Hellinger distance, we reach the conclusion of this part:
$H_{B}(\eta,\mathcal{P}_{k_{*}}(\Omega),h)\lesssim\log(1/\eta).$
Hence, the proof is completed.
K.2 Proof of Theorem J.2
@@
  • Inequality A. $\inf_{G\in\mathcal{G}_{k_{*}}(\Omega):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0$;
  • Inequality B. $\inf_{G\in\mathcal{G}_{k_{*}}(\Omega):\mathcal{D}_{1}(G,G_{*})>\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0$,
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{k_{*}}(\Omega):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon}\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0.$
Assume that the above inequality is not true, then, there exists a sequence of mixing measure $G_{n}:=\sum_{i=1}^{k_{*}}\exp(\beta^{n}_{i})\delta_{(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})}\in\mathcal{G}_{k_{*}}(\Omega)$ such that both $\mathcal{D}_{1}(G_{n},G_{*})$ and $\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{1}(G_{n},G_{*})$ go to zero as $n\to\infty$. Now, we define
@@
for any $j\in[k_{*}]$ as Voronoi cells with respect to the mixing measure $G_{n}$, where we denote $\theta^{n}_{i}:=(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})$ and $\theta^{*}_{j}:=(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$. In this proof, since our arguments are assymptotic, we can assume without loss of generality (WLOG) that these Voronoi cells does not depend on $n$, that is, $\mathcal{A}_{j}=\mathcal{A}^{n}_{j}$. Next, it follows from the hypothesis $\mathcal{D}_{1n}:=\mathcal{D}_{1}(G_{n},G_{*})\to 0$ as $n\to\infty$ that each Voronoi cell contains only one element. Therefore, we may assume WLOG that $\mathcal{A}_{j}=\{j\}$ for any $j\in[k_{*}]$, which implies that $(W_{j}^{n},a_{j}^{n},b_{j}^{n},\nu_{j}^{n})\to(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ and $\exp(\beta^{n}_{j})\to\exp(\beta_{j}^{*})$ as $n\to\infty$.
Subsequently, to specify the top $K$ selection in the formulations of $p_{G_{n}}(Y|X)$ and $p_{G_{*}}(Y|X)$, we divide the covariate space $\mathcal{X}$ into some subsets in two ways. In particular, we first consider $q:=\binom{k_{*}}{K}$ different $K$-element subsets of $[k_{*}]$, which are assumed to take the form $\{\tau_{1},\ldots,\tau_{K}\}$, for $\tau\in[q]$. Additionally, we denote $\{\tau_{K+1},\ldots,\tau_{k_{*}}\}:=[k_{*}]\setminus\{\tau_{1},\ldots,\tau_{K}\}$. Then, we define for each $\tau\in[q]$ two following subsets of $\mathcal{X}$:
$\mathcal{X}^{n}_{\tau}:=\Big{\{}x\in\mathcal{X}:-\|W_{j}-x\|\geq-\|W_{j^{\prime}}-x\|:\forall j\in\{\tau_{1},\ldots,\tau_{K}\},j^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\Big{\}},$
$\mathcal{X}^{*}_{\tau}:=\Big{\{}x\in\mathcal{X}:-\|W^{*}_{j}-x\|\geq-\|W^{*}_{j^{\prime}}-x\|:\forall j\in\{\tau_{1},\ldots,\tau_{K}\},j^{\prime}\in\{\tau_{K+1},\ldots,\tau_{k_{*}}\}\Big{\}}.$
Since $(W_{j}^{n},\beta^{n}_{j})\to(W_{j}^{*},\beta_{j}^{*})$ as $n\to\infty$ for any $j\in[k_{*}]$, we have for any arbitrarily small $\eta_{j}>0$ that $\|W_{j}^{n}-W_{j}^{*}\|\leq\eta_{j}$ and $|\beta^{n}_{j}-\beta_{j}^{*}|\leq\eta_{j}$ for sufficiently large $n$. By applying Lemma L.1, we obtain that $\mathcal{X}^{n}_{\tau}=\mathcal{X}^{*}_{\tau}$ for any $\tau\in[q]$ for sufficiently large $n$.
WLOG, we assume that
$\mathcal{D}_{1}(G_{n},G_{*})=\sum_{i=1}^{K}\Big{[}\exp(\beta_{i}^{n})\Big{(}\|\Delta W_{i}^{n}\|+\|\Delta a_{i}^{n}\|+\|\Delta b_{i}^{n}\|+\|\Delta\nu_{i}^{n}\|\Big{)}+\Big{|}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{|}\Big{]},$
@@
Let $\tau\in[q]$ such that $\{\tau_{1},\ldots,\tau_{K}\}=\{1,\ldots,K\}$. Then, for almost surely $(X,Y)\in\mathcal{X}^{*}_{\tau}\times\mathcal{Y}$, we can rewrite the conditional densities $p_{G_{n}}(Y|X)$ and $p_{G_{*}}(Y|X)$ as
$p_{G_{n}}(Y|X)$ $=\sum_{i=1}^{K}\frac{\exp(-\|W_{i}^{n}-X\|+\beta_{i}^{n})}{\sum_{j=1}^{K}\exp(-\|W_{j}^{n}-X\|+\beta^{n}_{j})}\cdot f(Y|(a_{i}^{n})^{\top}X+b_{i}^{n},\nu_{i}^{n}),$
$p_{G_{*}}(Y|X)$ $=\sum_{i=1}^{K}\frac{\exp(-\|W_{i}^{*}-X\|+\beta_{i}^{*})}{\sum_{j=1}^{K}\exp(-\|W_{j}^{*}-X\|+\beta_{j}^{*})}\cdot f(Y|(a_{i}^{*})^{\top}X+b_{i}^{*},\nu_{i}^{*}).$
@@
In this step, we aim to decompose the term $Q_{n}:=\Big{[}\sum_{i=1}^{K}\exp(-\|W_{i}^{*}-X\|+\beta_{i}^{*})\Big{]}\cdot[p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)]$, which can be represented as follows:
$Q_{n}$ $=\sum_{i=1}^{K}\exp(\beta_{i}^{n})\Big{[}F(Y|X;W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})-F(Y|X;W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})\Big{]}$
$-\sum_{i=1}^{K}\exp(\beta_{i}^{n})\Big{[}H(Y|X;W_{i}^{n})-H(Y|X;W_{i}^{*})\Big{]}$
$+\sum_{i=1}^{K}\Big{[}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{]}\Big{[}F(Y|X;W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})-H(Y|X,W_{i}^{*})\Big{]}$
@@
$A_{n}$ $=\sum_{i=1}^{K}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$=\sum_{i=1}^{K}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$A_{n}$ $=\sum_{i=1}^{K}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{2^{\alpha_{4}}\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$=\sum_{i=1}^{K}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\sum_{\begin{subarray}{c}\alpha_{3}+2\alpha_{4}=\eta,\\ 0\leq\alpha_{3}+\alpha_{4}\leq 1-|\alpha_{1}|-|\alpha_{2}|\end{subarray}}\frac{\exp(\beta_{i}^{n})}{2^{\alpha_{4}}\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$

### v4

@@ I.3 Over-specified Setting
  • Inequality A. $\inf_{G\in\mathcal{G}_{k_{*}}(\Theta):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0$;
  • Inequality B. $\inf_{G\in\mathcal{G}_{k_{*}}(\Theta):\mathcal{D}_{1}(G,G_{*})>\varepsilon^{\prime}}\dfrac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0$,
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{k_{*}}(\Theta):\mathcal{D}_{1}(G,G_{*})\leq\varepsilon}\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{\mathcal{D}_{1}(G,G_{*})}>0.$
Assume that the above inequality is not true, then, there exists a sequence of mixing measure $G_{n}:=\sum_{i=1}^{k_{*}}\exp(\beta^{n}_{i})\delta_{(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})}\in\mathcal{G}_{k_{*}}(\Theta)$ such that both $\mathcal{D}_{1}(G_{n},G_{*})$ and $\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]/\mathcal{D}_{1}(G_{n},G_{*})$ go to zero as $n\to\infty$. Now, we define
@@
for any $j\in[k_{*}]$ as Voronoi cells with respect to the mixing measure $G_{n}$, where we denote $\theta^{n}_{i}:=(W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})$ and $\theta^{*}_{j}:=(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$. In this proof, since our arguments are assymptotic, we can assume without loss of generality (WLOG) that these Voronoi cells does not depend on $n$, that is, $\mathcal{A}_{j}=\mathcal{A}^{n}_{j}$. Next, it follows from the hypothesis $\mathcal{D}_{1n}:=\mathcal{D}_{1}(G_{n},G_{*})\to 0$ as $n\to\infty$ that each Voronoi cell contains only one element. Therefore, we may assume WLOG that $\mathcal{A}_{j}=\{j\}$ for any $j\in[k_{*}]$, which implies that $(W_{j}^{n},a_{j}^{n},b_{j}^{n},\nu_{j}^{n})\to(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ and $\exp(\beta^{n}_{j})\to\exp(\beta_{j}^{*})$ as $n\to\infty$. Then, the loss function between $G_{n}$ and $G_{*}$ is given by
$\mathcal{D}_{1}(G_{n},G_{*})=\sum_{i=1}^{k_{*}}\Big{[}\exp(\beta_{i}^{n})\Big{(}\|\Delta W_{i}^{n}\|+\|\Delta a_{i}^{n}\|+\|\Delta b_{i}^{n}\|+\|\Delta\nu_{i}^{n}\|\Big{)}+\Big{|}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{|}\Big{]},$
@@
In this step, we aim to decompose the term $Q_{n}:=\Big{[}\sum_{i=1}^{k_{*}}\exp(-\|W_{i}^{*}-X\|+\beta_{i}^{*})\Big{]}\cdot[p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)]$, which can be represented as follows:
$Q_{n}$ $=\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\Big{[}F(Y|X;W_{i}^{n},a_{i}^{n},b_{i}^{n},\nu_{i}^{n})-F(Y|X;W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})\Big{]}$
$-\sum_{i=1}^{k_{*}}\exp(\beta_{i}^{n})\Big{[}H(Y|X;W_{i}^{n})-H(Y|X;W_{i}^{*})\Big{]}$
$+\sum_{i=1}^{k_{*}}\Big{[}\exp(\beta_{i}^{n})-\exp(\beta_{i}^{*})\Big{]}\Big{[}F(Y|X;W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*})-H(Y|X,W_{i}^{*})\Big{]}$
@@
$A_{n}$ $=\sum_{i=1}^{k_{*}}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$=\sum_{i=1}^{k_{*}}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$A_{n}$ $=\sum_{i=1}^{k_{*}}\sum_{|\alpha|=1}\frac{\exp(\beta_{i}^{n})}{2^{\alpha_{4}}\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$
@@
$=\sum_{i=1}^{k_{*}}\sum_{|\alpha_{1}|=0}^{1}\sum_{|\alpha_{2}|=0}^{1-|\alpha_{1}|}\sum_{\eta=0}^{2(1-|\alpha_{1}|-|\alpha_{2}|)}\sum_{\begin{subarray}{c}\alpha_{3}+2\alpha_{4}=\eta,\\ 0\leq\alpha_{3}+\alpha_{4}\leq 1-|\alpha_{1}|-|\alpha_{2}|\end{subarray}}\frac{\exp(\beta_{i}^{n})}{2^{\alpha_{4}}\alpha!}\cdot(\Delta W_{i}^{n})^{\alpha_{1}}(\Delta a_{i}^{n})^{\alpha_{2}}(\Delta b_{i}^{n})^{\alpha_{3}}(\Delta\nu_{i}^{n})^{\alpha_{4}}$